7. A large stone is resting on the bottom of the swimming pool. The normal force of the bottom of the pool on the stone is equal to the a. weight of the stone. b. weight of the water displaced. c. difference between the weight of the stone and the weight of the displaced water. d. sum of the weight of the stone and the weight of the displaced water.
8. Legend says that Archimedes, in determining whether or not the kings crown was made of pure gold, measured its volume by the water displacement method.If the density of gold is 19.3 /cm3, and the crowns mass is 600 g, what volume would be necessary to prove that it is pure gold? a. 3.1E+1 cm b. 3.2E-2 cm c. 3.2E+1 cm d. 3,0E+1 cm e. 1.2E+4 cm
Answer #1

7)correct ans:

(c)difference between the weight of the stone and the weight of water

-----------------

8)Correct Ans:

(a)3.1 E+1 cm^3

We know that

d = m/v => V = m/d

V = 600/19.3 = 31 cm^3

V = 3.1 E+1 cm^3